OpenStudy (cooltowl):

How many chlorine atoms are present in 25.3 g of the compound?

OpenStudy (cooltowl):

@Shalante

OpenStudy (cooltowl):

@matt101

OpenStudy (matt101):

What compound are you talking about?

OpenStudy (cooltowl):

CCl4

OpenStudy (matt101):

We want to know how many Cl atoms there are. When we're looking for a number, always convert to moles! We have the mass of CCl4, and we can find moles by dividing the mass by the molar mass of CCl4, which is 154 g/mol. moles of CCl4 = 25.3 g / 154 g/mol = 0.1643 mol Each molecules of CCl4 contains 4 atoms of Cl, so for however many moles of CCl4 we have, there are FOUR times as many moles of Cl - in this case, 0.1643 x 4 = 0.6572 mol. Now we want to know the number of ATOMS of Cl. We can go from moles to number of atoms by multiplying moles by Avogadro's number: 6.022 x 10^23 atoms of Cl = 0.6572 mol x 6.022 x 10^23 = 3.958 x 10^23 atoms Does that make sense?

OpenStudy (cooltowl):

That makes perfect sense. Thank you so much for the step by step explanation:)